OpenStudy (anonymous):

If A,B,C are the angles of a triangle then \[\left[\begin{matrix}\cos^2A & CotA & 1 \\ \cos^2B & cotB & 1 \\ \cos^2C & cotC & 1\end{matrix}\right]\] Find Its Determinant ?

OpenStudy (reemii):

what about it?

OpenStudy (anonymous):

Oh..No....Sorry We Have to Find Its Determinant

OpenStudy (reemii):

cot A = cos(A)/sin(A) ?

OpenStudy (anonymous):

I did Everything but they all Luk Werid.. :(

OpenStudy (reemii):

did you try \(L_1\leftarrow L_1-L_3\) \(L_2 \leftarrow L_2-L_3\) ?

OpenStudy (anonymous):

the last column is a linear transform of the first column. \[ \left[\begin{matrix}\cos^2A&\cot A &\cos^2A\\ \cos^2B&\cot B &\cos^2B\\ \cos^2C&\cot C &\cos^2C\\ \end{matrix}\right]+ \left[\begin{matrix} 0&0&\sin^2A\\ 0&0&\sin^2B\\ 0&0&\sin^2C\\ \end{matrix}\right] \]

OpenStudy (reemii):

but det(A+B) is not something as easy as det(A)+det(B). What I meant is: \(\left[\begin{array}{ccc} \cos^2A-\cos^2C & \cot A-\cot C & 0\\ \cos^2B-\cos^2C & \cot B-\cot C & 0\\ \cos^C & \cot C & 1 \end{array}\right] \) so the determinant is \[ (\cos^2A-\cos^2C)(\cot B-\cot C) - (\cot A-\cot C)(\cos^2B-\cos^2C)\\ =\cos^2A\cot B - \cos^2 A\cot C - \cos^2C\cot B\\ \:\: -(\cot A\cos^2 - \cot A\cos^2C - \cot C\cos^2B) \]